# The resolution quiver of Nakayama algebras which are minimal Auslander-Gorenstein

### v1

@@ 4. Proof of the main result
In this section, we study the dominant dimensions of Nakayama algebras. Then we prove the main result of the present paper.
@@
of $M$. The codominant dimension $\mathrm{codom}.\dim_{A}M$ of $M$ is the maximal integer $n$ or $\infty$ such that $P_{j}$ is injective for all $j<n$. The left dominant dimension of $A$ is equal to the right dominant dimension of $A$; see [8]. It is known that any Nakayama algebra has dominant dimension at least $1$.
For an integer $d\geq 1$, an algebra $A$ is called higher Auslander if
@@
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $1$ if and only if
@@
    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
@@
$\implies$" Since $\mathrm{id}_{A}\,A=1$, we have $\mathrm{gl}.\dim A=1$ by Lemma 2.4. Then by Lemma 2.3 $R(A)$ is connected of weight $1$. There is a direct sum decomposition
$A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r}$
where $A_{i}$ is a hereditary linear Nakayama algebra for all $i\leq r$. Let $S$ be a cyclic vertex in $R(A)$. Then $S$ is the simple module at the source of $A_{i}$ and $\tau^{-1}{S}$ is the simple module at the sink of $A_{i-1}$. Since $\tau^{-1}S$ is projective, it is black.
@@
$0\to\Delta(S)\to P(\tau T)\to P(T)\to T\to 0$
such that $\tau^{-1}S$ is black. Then $P(S)$ is injective by Lemma 2.1. Since $\mathrm{gl}.\dim A$ is at most $2$, by (3.2) we have $\Delta(S)=P(S)$. Then the sequence splits and $T$ is projective. We conclude that the selfinjective dimension of $A$ is $1$. ∎
@@
Let $A$ be a Nakayama algebra of global dimension $1$. Since not all components of $A$ are simple, there is at least one red cyclic vertex in $R(A)$.
@@
Let $A$ be a Nakayama algebra with Kupisch series $(1,3,2,1)$. Then $A$ is the direct sum of two hereditary linear Nakayama algebras. The resolution quiver of $R(A)$ is the following. Observe that the cyclic vertex $2$ is a red vertex.
@@
Let $A$ be a Nakayama algebra of dominant dimension at least $2$ and let $S$ be a simple $A$-module. If $\mathrm{pd}_{A}\,\tau^{-1}S=1$, then $\mathrm{id}_{A}\,S\neq 1$.
@@
Since $\mathrm{pd}_{A}\,\tau^{-1}S=1$, then $\tau^{-1}S$ is red. Let $m$ be the minimal integer such that $\tau^{-m-1}S$ is black. By Lemma 2.1 we know that $\gamma S=\gamma\tau^{-m}S$ and $P(\tau^{-m}S)$ is injective. Consider the exact sequence
$0\to P(S)\to P(\tau^{-m}S)\to I(\tau^{-1}S).$
Since $\mathrm{dom}.\dim A\geq 2$, the module $I(\tau^{-1}S)$ is projective. Then the injective dimension of $S$ is not equal to $1$. ∎
We have the following.
@@
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$ if and only if
@@
(2) Suppose three distinct vertices $T,\tau T,\tau^{2}T$ have the same cyclic successor in $R(A)$. Then $T$ and $\tau T$ are red. Since $\mathrm{dom}.\dim A\geq 2$, by Lemma 4.4 $\tau T$ and $\tau^{2}T$ are non-leaf vertices. By (1) we get that $\tau T,\tau^{2}T$ are cyclic. This is impossible.
(3) Let $S$ be a red vertex in $R(A)$. By (2) $\tau^{-1}S$ is black. By Lemma 2.1 $P(S)$ is injective. Since the selfinjective dimension of $A$ is $2$, there is an exact sequence
@@
Then the injective dimension of $S$ is equal to $1$. By Lemma 2.5 the vertex $S$ is a leaf in $R(A)$. In particular, $S$ is not cyclic.
@@
By Remark 4.2 the algebra $A$ is not hereditary. Then the selfinjective dimension of $A$ is $2$. It remains to prove $\mathrm{dom}.\dim A\geq 2$. By [9, 6.2], it suffices to show that each red simple $A$-module is in the image of black simple $A$-modules under $\tau^{-1}\gamma\tau$.
Since $\operatorname{dist}A=1$ and all cyclic vertices are black. We infer that all black vertices are cyclic. Let $S$ be a red vertex in $R(A)$. Since $\gamma(S)$ has at most two predecessors, $\tau(S)$ is black and then cyclic. There is a cyclic vertex $T$ in $R(A)$ such that $\gamma(T)=\tau(S)$. If $\tau^{-1}T$ is black, then
$\tau^{-1}\gamma\tau(\tau^{-1}T)=\tau^{-1}\gamma T=\tau^{-1}\tau S=S.$
If $\tau^{-1}T$ is red, then $\tau^{-2}T$ is black. We have $\gamma\tau^{-1}T=\gamma T$ and
$\tau^{-1}\gamma\tau(\tau^{-2}T)=\tau^{-1}\gamma\tau^{-1}T=\tau^{-1}\gamma T=\tau^{-1}\tau S=S.$
This finishes our proof. ∎
@@
Let $A$ be a Nakayama algebra with Kupisch series $(6,7,6,7)$. Then $R(A)$ is the following.
@@
It is easy to see that $R(A)$ fulfills the condition in Proposition 4.5. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$.
For algebras of dominant dimension at least $3$, we need the following key result; see [15, 4.1 and 4.2]. We give a short proof here.
@@
    $\tau^{-1}S$ and $S$ are black for every leaf $S$ in $R(A)$;
@@
$(1)\implies(2)$ Let $S$ be a leaf in $R(A)$. Since $\mathrm{dom}.\dim A\geq 2$, we know from Lemma 4.4 that $\tau^{-1}S$ is black. Then $P(S)$ is injective by Lemma 2.1. If $S$ is red, there is an exact sequence
@@
Since $\mathrm{dom}.\dim A\geq 3$, the module $I(\tau^{-1}S)$ is projective and the sequence splits. This is impossible. Then $S$ is black.
@@
$(3)\iff(4)$ Since $R(A)^{\prime}$ and $R(\varepsilon{A})$ are isomorphic quivers, the defect of $\varepsilon(A)$ is equal to the number of leaf vertices in $R(A)^{\prime}$. Then $\operatorname{def}A$ is equal to $\operatorname{def}\varepsilon(A)$ if and only if $S$ is the unique predecessor of its successor for every leaf $S$ in $R(A)$.
$(4)\implies(5)$ Let $T$ be a leaf in $R(A)$ and $S$ be the successor of $T$. Since $\tau^{-1}T$ is a black leaf, by (3.3) there is a non-split exact sequence
$0\to J(\tau^{-1}S)\to P(T)\to P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0.$
Since $A$ satisfies (B), by Proposition 3.9 the projective injective objects in $\mathcal{F}$ and the projective injective $A$-modules contained in $\mathcal{F}$ coincide. Then
@@
Note that dominant dimension at least $3$ implies (B) and (B) for any Nakayama algebra. We have the following.
@@
Let $A$ be a Nakayama algebra of dominant dimension at least $3$ and let $\Delta(S)$ be a vertex in $R(\varepsilon{A})$. Then
@@
    $\tau^{-1}_{\varepsilon}\Delta(S)$ is black if and only if $\tau^{-1}S$ is black.
@@
This follows from Corollary 3.6 and Lemma 4.7. Compare [15, 4.4.2]. ∎
Now we restate and prove Theorem A. It gives a criterion to decide whether a Nakayama algebra is minimal Auslander-Gorenstein of odd selfinjective dimension.
@@
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. For any $m\geq 1$, the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1 if and only if
@@
    $\tau^{-1}$ of all cyclic vertices in $R(A)$ are black.
@@
$\implies$" We prove by induction on $m$. If $m=1$, this is proven in Proposition 4.1. If $m\geq 2$, by Corollary 4.9 the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3. By induction hypothesis, the following statements hold for $\varepsilon(A)$.
@@
    $\tau^{-1}_{\varepsilon}$ of all cyclic vertices in $R(\varepsilon A)$ are black.
There is an injective map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$ which preserves the weight. Then (1), (2) and (3) hold. By Corollary 4.8 we get (4).
@@
There is an injective map of resolution quivers $\varphi$ from $R(\varepsilon{A})$ to $R(A)$ which preserves the weight. Then (1a), (2a) and (3a) hold. By Corollary 4.8 we obtain (4a).
By the induction hypothesis, the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3. Since the dominant dimension of $A$ is at least $3$, by Corollary 4.9 the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1. ∎
We restate Theorem B as follows. The proof is similar, we omit it here.
@@
Let $A$ be a Nakayama algebra and $R(A)$ be its resolution quiver. For any $m\geq 1$, the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2m$ if and only if
@@
For a Nakayama algebra $A$, denote by $\operatorname{cyc}A$ be the number of cyclic vertices in its resolution quiver. We have the following.
@@
Let $A$ be a Nakayama algebra. Then $\operatorname{def}A\leq\operatorname{cyc}A$ provided that $A$ is minimal Auslander-Gorenstein of even selfinjective dimension.
@@ 5. Applications and examples
Let $A$ be a Nakayama algebra of rank $n$ with Kupisch series
@@
Let $A$ be a Nakayama algebra and let $d$ be an even number. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $d$ if and only if so is $\bar{A}$.
For any integer $n$, denote by $\mathbb{A}_{n}$ be the hereditary linear Nakayama algebra of rank $n$. For any $m\leq n$, denote by $N_{n}(m)$ be the quotient algebra of $\mathbb{A}_{n}$ via the ideal generated by all paths of length $m$. The Kupisch series of $N_{n}(m)$ is
@@
The algebra $N_{n}(m)$ has finite global dimension. If $m=2$, then $N_{n}(m)$ is higher Auslander of global dimension n1n-1.
Suppose $m>2$ and $N_{n}(m)$ is a higher Auslander algebra. Write
@@
such that $q,r$ are integers and $0\leq r\leq m-1$. Since $S(1)$ is cyclic and $\gamma^{-1}S(1)$ consists of $m$ simple $A$-modules, the global dimension of $N(n,r)$ is odd by Theorem 4.11. Since S(nr+1)S(n-r+1) is cyclic, then S(nr)S(n-r) is black by Theorem 4.10. We infer that $r=0$ and $m$ divides $n$.
@@
The algebra $N_{n}(m)$ is higher Auslander if and only if $m=2$ or $m$ divides $n$. In both cases, the global dimension of $N_{n}(m)$ is 2n/m12n/m-1.
Following [15], for any Nakayama algebra $A$, there is a unique Nakayama algebra $\varepsilon^{-1}{A}$ of dominant dimension at least $3$ such that $\varepsilon(\varepsilon^{-1}{A})$ is isomorphic to $A$. Let $A$ be a Nakayama algebra of rank $n$ with Kupisch series
@@
Denote by $d_{a}$ the maximum of $c_{a}-c_{a-1}$ and 0 for each $a\leq n$, and denote by $d$ the sum of all $d_{a}$. Let $\tilde{c}_{a}$ be a series of length $1+d_{a+1}$ such that
$(\tilde{c}_{a})_{\ell}=c_{a}+\sum_{i=1}^{c_{a}+\ell-1}d_{a+i}$
for each $\ell\leq 1+d_{a+1}$. Then $\varepsilon^{-1}{A}$ is a Nakayama algebra of rank $n+d$ with Kupisch series
@@
One sees that $\chi(\varepsilon^{-1}{A})$ is equal to the number of simple components of $A$.
Recall the following result in [15].
@@
Let $d\geq 3$ and $\Lambda$ be a cyclic Nakayama algebra. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
@@
where $A$ is a higher Auslander cyclic Nakayama algebra of global dimension d2d-2 or
@@
where $A_{i}$ is a higher Auslander linear Nakayama algebra of global dimension d2d-2 for $i\leq r$.
Similarly, we have the following.
@@
Let $d\geq 3$ and $\Lambda$ be a linear Nakayama algebra. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
@@
where $A_{i}$ is a higher Auslander linear Nakayama algebra of global dimension d2d-2 for $i\leq r$.
Connected Nakayama algebras of which are higher Auslander of global dimension $1$ are hereditary. Connected Nakayama algebras of which are higher Auslander of global dimension $2$ are just the Auslander algebras of radical square zero connected Nakayama algebras. We can repeatedly apply Proposition 5.4 and 5.3 to give all connected Nakayama algebras of which are higher Auslander.
@@
The Kupisch series of $\varepsilon^{-1}(A)$ is
@@
Then $\varepsilon^{-1}(A)$ is higher Auslander of global dimension $3$.

### v2

@@ 4. Proof of the main result
In this section, we study the dominant dimension of Nakayama algebras by means of resolution quivers. Then we prove the main result of the present paper.
@@
of $M$. The codominant dimension $\mathrm{codom}.\dim_{A}M$ of $M$ is the maximal integer $n$ or $\infty$ such that $P_{j}$ is injective for all $j<n$. The left and right dominant dimensions of $A$ are equal; see [8]. It is known that the dominant dimension of any Nakayama algebra is at least $1$.
Following [4, 5], for an integer $d\geq 1$, an algebra $A$ is called higher Auslander if
@@
Higher Auslander algebras are just the minimal Auslander-Gorenstein algebras of finite global dimension.
@@
Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then the selfinjective dimension of $A$ is equal to $1$ if and only if
@@
    all cyclic vertices in $R(A)$ are *-black.
@@
$\implies$" Since $\mathrm{id}_{A}\,A=1$, by Lemma 2.4 we have $\mathrm{gl}.\dim A=1$. There is a direct sum decomposition of algebras
$A=A_{1}\oplus A_{2}\oplus\cdots\oplus A_{r},$
where $A_{i}$ is a hereditary linear Nakayama algebra for each $i\leq r$. By Lemma 2.3 $\operatorname{dist}A=1$ and $R(A)$ is connected of weight $1$. Let $S$ be a cyclic vertex in $R(A)$. Then $S$ is the simple module at the source of $A_{i}$ and $\tau^{-1}{S}$ is the simple module at the sink of $A_{i-1}$. Since $\tau^{-1}S$ is projective, it is black.
@@
$0\to\Delta(S)\to P(\tau T)\overset{\alpha}{\to}P(T)\to T\to 0.$
Since $R(A)$ is connected of weight $1$, by Lemma 2.3 the global dimension of $A$ is at most $2$. It follows from (3.2) that $\Delta(S)=P(S)$. Then $\alpha$ is zero and $T$ is projective. We conclude that the selfinjective dimension of $A$ is equal to $1$. ∎
@@
Let $A$ be a Nakayama algebra of global dimension $1$. Since not all connected component of $A$ are simple, there exists at least one black leaf in $R(A)$.
@@
Let $A$ be a Nakayama algebra with Kupisch series $(1,3,2,1)$. Then $A$ is the direct sum of two hereditary linear Nakayama algebras. The resolution quiver of $R(A)$ is the following.
@@
Observe that $4$ is a black leaf in the resolution quiver of $A$.
@@
Let $A$ be a Nakayama algebra. The following are equivalent.
  1. (1)
    $\mathrm{dom}.\dim A\geq 2$;
  2. (2)
    every leaf in $R(A)$ is *-black;
  3. (3)
    every *-red vertex in $R(A)$ is the successor of a *-black vertex.
@@
$(1)\implies(2)$ Let $S$ be a *-red vertex in $R(A)$. Take a *-black vertex $T$ such that $\gamma(T)=\gamma(S)$. Then $P(T)$ is injective. There is an exact sequence
$0\to P(S)\to P(T)\to I(\tau^{-1}S).$
Since $\mathrm{dom}.\dim A\geq 2$, we have $I(\tau^{-1}S)$ is projective. Then $S$ is not a leaf.
$(2)\implies(3)$ Let $S$ be a *-red vertex in $R(A)$. Since $S$ is not a leaf, there is a *-black vertex $T$ such that $\gamma(T)=S$. Then $S$ is the successor of a *-black vertex.
$(3)\implies(1)$ This follows from [9, 6.2]. Let $P(T)$ be an indecomposable non-injective projective $A$-module. Then $T$ is *-red and there is a *-black simple $A$-module $S$ such that $\gamma(S)=T$. Since $S$ is *-black, $P(S)$ and $I(\tau^{-1}T)$ are isomorphic. Then we have $\mathrm{dom}.\dim A\geq 2$. ∎
@@
Let $A$ be a Nakayama algebra and $R(A)$ its resolution quiver. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$ if and only if
@@
(2) Suppose three distinct vertices $T,\tau T,\tau^{2}T$ have the same successor in $R(A)$. Then $T$ and $\tau T$ are red. Since $\mathrm{dom}.\dim A\geq 2$, by Lemma 4.4 $\tau T$ and $\tau^{2}T$ are non-leaf vertices. Since $\operatorname{dist}A=1$, both $\tau T$ and $\tau^{2}T$ are cyclic. This is impossible.
(3) Let $S$ be a red vertex in $R(A)$. By Lemma 4.4 $\tau^{-1}S$ is black. Then $P(S)$ is injective. Since $\mathrm{id}_{A}\,A=2$, there is an exact sequence
@@
By Lemma 2.5 $S$ is a leaf in $R(A)$. In particular, $S$ is not cyclic.
@@
Let $S$ be a leaf vertex in $R(A)$. Since $\operatorname{dist}A=1$ and all cyclic vertices are black, then $S$ is red and $\gamma(S)=\gamma(\tau S)$. Since $\gamma(S)$ has at most $2$ predecessors, $\tau^{-1}S$ is black. By Lemma 4.4, the dominant dimension of $A$ is at least $2$. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2$. ∎
@@
Let $A$ be a Nakayama algebra with Kupisch series $(6,7,6,7)$. The resolution quiver $R(A)$ is the following.
@@
It is easy to check that $A$ fulfills the conditions in Proposition 4.5. Then $A$ is a minimal Auslander-Gorenstein algebra of selfinjective dimension $2$.
For algebras of dominant dimension at least $3$, recall the following key result from [16, 4]. We give a short proof here.
@@
    every leaf in $R(A)$ is black and *-black;
@@
$(1)\implies(2)$ Let $S$ be a leaf in $R(A)$. Since $\mathrm{dom}.\dim A\geq 2$, by Lemma 4.4 $S$ is *-black and $P(S)$ is injective. Assume $S$ is red. There is an exact sequence
@@
Since $\mathrm{dom}.\dim A\geq 3$, then $I(\tau^{-1}S)$ is projective. This is impossible. We conclude that $S$ is black and *-black.
@@
$(3)\iff(4)$ Note that $\operatorname{def}\varepsilon(A)$ is equal to the number of leaf vertices in $R(A)^{\prime}$. Then $\operatorname{def}A$ and $\operatorname{def}\varepsilon(A)$ are equal if and only if every leaf in $R(A)$ is the unique predecessor of its successor.
$(2)\implies(5)$ Let $T$ be a leaf in $R(A)$ and $S$ the successor of $T$. Since $\tau^{-1}T$ is black, by (3.3) there is an exact sequence
$0\to J(\tau^{-1}S)\to P(T)\overset{w}{\to}P(I(\tau^{-1}T))\to I(\tau^{-1}T)\to 0.$
Since $T$ is a leaf, we have $w\neq 0$. Note that $A$ satisfies (B). By Proposition 3.5 any projective injective object in $\mathcal{F}$ is a projective injective $A$-module. Then
@@
Note that $\mathrm{dom}.\dim A\geq 3$ implies (B) and (B). We have the following.
@@
Let $A$ be a Nakayama algebra of dominant dimension at least $3$ and $\Delta(S)$ a vertex in $R(\varepsilon{A})$. Then
@@
    $\Delta(S)$ is *-black if and only if $S$ is *-black.
Combining Lemma 4.4 and 4.7 with the previous result, we have the following.
@@
Let $A$ be a Nakayama algebra and $m\geq 0$.
  1. (1)
    The following are equivalent.
    1. (a)
      $\mathrm{dom}.\dim A\geq 2m+1$;
    2. (b)
      $S,\gamma(S),\cdots,\gamma^{m-1}(S)$ are black and *-black for every leaf $S$;
    3. (c)
      $|\gamma^{-1}\gamma^{i}(S)|=1$ for every leaf $S$ and $i\leq m$.
  2. (2)
    The following are equivalent.
    1. (a)
      $\mathrm{dom}.\dim A\geq 2m+2$;
    2. (b)
      $S,\gamma(S),\cdots,\gamma^{m-1}(S)$ are black and *-black, and $\gamma^{m}(S)$ is *-black for every leaf $S$;
    3. (c)
      $|\gamma^{-1}\gamma^{m}(S)|=1$ and $\gamma^{m}(S)$ is *-black for every leaf $S$ and $i\leq m$.
We have the following; compare [16, 4.4.2].
Proposition 4.10.
@@
This follows from Lemma 4.7 and Corollary 3.6. ∎
Now we restate and prove Theorem A.
@@
Let $A$ be a Nakayama algebra and $m\geq 1$. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1 if and only if
@@
    all cyclic vertices in $R(A)$ are *-black.
@@
$\implies$" We prove by induction on $m$. If $m=1$, this is done in Proposition 4.1. Suppose $m\geq 2$. By Proposition 4.10 the syzygy filtered algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3. By induction hypothesis, the following statements hold.
@@
    all cyclic vertices in $R(\varepsilon A)$ are *-black.
Since there is an injective weight preserving map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$, then (1), (2) and (3) hold. By Corollary 4.8 the statement (4) holds.
@@
Since there is an injective weight preserving map of resolution quivers from $R(\varepsilon{A})$ to $R(A)$, then (1a), (2a) and (3a) hold. By Corollary 4.8 the statement (4a) holds.
By the induction hypothesis, the algebra $\varepsilon(A)$ is minimal Auslander-Gorenstein of selfinjective dimension 2m32m-3. Since the dominant dimension of $A$ is at least $3$, by Proposition 4.10 the algebra $A$ is minimal Auslander-Gorenstein of selfinjective dimension 2m12m-1. ∎
Let us restate Theorem B as follows. The proof is similar, we omit it here.
@@
Let $A$ be a Nakayama algebra and $m\geq 1$. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $2m$ if and only if
@@
For a Nakayama algebra $A$, denote by $\operatorname{cyc}A$ the number of cyclic vertices in its resolution quiver.
@@
Let $A$ be a Nakayama algebra. If $A$ is a minimal Auslander-Gorenstein algebra of even selfinjective dimension, then $\operatorname{cyc}A\leq\operatorname{def}A$.
@@ 5. Applications and examples
In this section, we present some applications and examples of our results.
Let $A$ be a Nakayama algebra with Kupisch series
@@
Let $A$ be a Nakayama algebra and $d$ an even number. Then $A$ is minimal Auslander-Gorenstein of selfinjective dimension $d$ if and only if so is $\bar{A}$.
Let $2\leq m\leq n$. Denote by $\mathbb{A}_{n}$ the hereditary linear Nakayama algebra of rank $n$. Let $N_{n}(m)$ be the quotient algebra of $\mathbb{A}_{n}$ via the ideal generated by all paths of length $m$. The algebra $N_{n}(m)$ is a linear Nakayama algebra and has finite global dimension. The Kupisch series of $N_{n}(m)$ is
@@
If $m=2$, then $N_{n}(m)$ is higher Auslander of global dimension n1n-1. Suppose $m>2$ and $N_{n}(m)$ is a higher Auslander algebra. We write
@@
such that $q,r$ are integers and $0\leq r\leq m-1$. Since $S(1)$ is cyclic and $S(1)$ has $m$ predecessors in the resolution quiver, the global dimension of $N_{n}(m)$ is odd by Theorem 4.12. Since S(nr+1)S(n-r+1) is cyclic, S(nr)S(n-r) is black by Theorem 4.11. It follows that $r=0$ and $m$ divides $n$.
@@
The algebra $N_{n}(m)$ is a higher Auslander algebra if and only if $m=2$ or $m$ divides $n$. In both cases, the global dimension of $N_{n}(m)$ is 2n/m12n/m-1.
For any Nakayama algebra $A$, by [16] there is a unique Nakayama algebra $\varepsilon^{-1}{A}$ of dominant dimension $\geq 3$ such that $\varepsilon(\varepsilon^{-1}{A})$ is naturally isomorphic to $A$. The algebra $\varepsilon^{-1}A$ is given as follows. Suppose the Kupisch series of $A$ is
@@
Denote by $d_{a}$ the maximum of $c_{a}-c_{a-1}$ and 0 for each $a\leq n$, and by $d$ the sum of all $d_{a}$. The subscript is taken modulo $n$. Let $\tilde{c}_{a}$ be a series of length $1+d_{a+1}$ with
[…]
Then $\varepsilon^{-1}{A}$ is a Nakayama algebra of rank $n+d$ with Kupisch series
@@
Note that $\chi(\varepsilon^{-1}A)$ equals the number of simple components of $A$.
Recall the following result; see [14, 16].
@@
Let $\Lambda$ be a cyclic Nakayama algebra and $d\geq 3$. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
@@
where $A$ is a cyclic Nakayama algebra which is higher Auslander of global dimension d2d-2 or isomorphic to
@@
where $A_{i}$ is a linear Nakayama algebra which is higher Auslander of global dimension d2d-2 for each $i\leq r$.
By Proposition 4.10, we have the following.
@@
Let $\Lambda$ be a linear Nakayama algebra and $d\geq 3$. Then $\Lambda$ is higher Auslander of global dimension $d$ if and only if $\Lambda$ is isomorphic to
@@
where $A_{i}$ is a linear Nakayama algebra which is higher Auslander of global dimension d2d-2 for each $i\leq r$.
@@
Then $\varepsilon^{-1}(A)$ is higher Auslander of global dimension $3$ with Kupisch series
@@
Nakayama algebras of global dimension $1$ are hereditary, and Nakayama algebras which are higher Auslander of global dimension $2$ are the Auslander algebras of Nakayama algebras with radical square zero. For any $d\geq 3$, we can repeatedly apply Proposition 5.4 and 5.3 to get any connected Nakayama algebra which is higher Auslander of global dimension $d$.
Example 5.6.
Let $n\geq m\geq 0$ and $A$ be a non-simple linear Nakayama algebra. The algebra
$\Lambda=\varepsilon^{-n}(\mathbb{A}_{1}^{m}\oplus A)$
is a connected Nakayama algebra; it is linear if and only if $n=m$. We have
  1. (1)
    $\varepsilon^{-m}(\mathbb{A}_{1}^{m}\oplus\mathbb{A}_{n})\cong N_{mn+n}(n)$ for any $m\geq 1$ and $n\geq 2$;
  2. (2)
    $\varepsilon^{-m}(\mathbb{A}_{1}^{m}\oplus N_{3}(2))\cong N_{2m+3}(2)$ for any $m\geq 1$.